Complete the following fission reaction. 239 94 Puton → + 91 38 Sr+35n 146 Ba 56 143 Xe 54 91 Sr 38 148 Ba 56 91 38 Mn
Answer #1

this question can be easily solved by balancing the mass number and atomic number of reactants and products of this reaction,

here reactants are Pu and a neutron and products: unknown element,Sr, 3 neutrons

total mass number of reactants= 239+1 =240 (a nuetron has mass number of 1)

let mass number of unknown element be x

total mass number of products=x+91+(3*1)=x+94

we know that ;total mass number of reactants=total mass number of products

so x+94=240; implies x=240-94=146 ;so mass number of unknown element =146

let atomic number of unknown element be y

total proton number of products=y+38+(3*0)=y+38

total proton number of reactants=94+0=94

total proton number of products=total proton number of reactants

so y+38=94 ; y=94-38=56 atomic number of unknown element is 56

now looking at options

146Ba56 matches our findings

so 146Ba56 is the answer

this problem is done based on conservation of mass.

thank you